Sam emptied his collection of nickels and dimes on his bed. If his 46 coins were worth $3.10,how many coins of each type did he have?
Answer by fractalier(2101)

Here n are nickels and d are dimes...so we have
n + d = 46
and
5n + 10d = 310 (working in cents now)
Multiply the first eqn by 5 and subtract, we get
5n + 10d = 310
-(5n + 5d = 230)
and we have
5d = 80
d = 16
n = 30